### 15 Apr 2026 v1

# Role of volatility mixing in wealth condensation transition

Abstract
We study the role of heterogeneous volatility in a networked wealth dynamics model and its impact on the wealth condensation transition. Extending the Bouchaud–Mézard framework, we introduce binary volatility in networks and investigate how its configuration affects the effective power-law tail exponent of the wealth distribution. Using a stochastic block model, we control the mixing between volatility groups and show that the effective exponent is governed not only by the global parameter $\Lambda=2J/\beta^{2}$ but also by the volatility configuration in the network. We find that local interactions between nodes with different volatility induce a neutralization of group-wise exponents, which lowers the aggregate tail exponent and can drive a condensation transition across $\gamma_{\rm c}=2$. Our results identify volatility mixing as another control mechanism for wealth condensation and highlight the importance of noise heterogeneity in nonequilibrium systems on networks.
I Introduction
Heavy-tailed wealth distributions are a hallmark of socioeconomic inequality, and the tail exponent $\gamma$ plays a central role in determining the degree of concentration. In the context of wealth dynamics, Bouchaud and Mézard [6] (BM) established a paradigmatic framework showing that the stationary wealth distribution exhibits a power-law tail whose exponent $\gamma$ depends only on the ratio between interaction strength and squared volatility, $\Lambda\equiv 2J/\beta^{2}$ (see Fig. 1), both in the mean-field (MF) case and Erdős–Rényi (ER) networks [9]. This framework suggests that the condensation transition, signaled by $\gamma\leq 2$, is governed by a single control parameter.
However, real systems are rarely homogeneous: individuals may differ not only in growth rates but also in volatility. Although heterogeneity in growth has been studied [3, 11], the effect of heterogeneous volatility remains largely unexplored. This is important because volatility represents intrinsic fluctuations in wealth and may interact with the network structure in ways that cannot be reduced to a single effective parameter.
In this paper, we study how volatility configuration affects wealth distribution and condensation in the BM model. Using a stochastic block model (SBM), we tune the mixing between high- and low-volatility groups while keeping degree statistics nearly unchanged. This allows us to isolate the volatility configuration effect itself.
We show that the effective tail exponent is determined not only by $\Lambda$ but also by how volatility is allocated on the network. In particular, mixing between nodes with different volatility induces a neutralization of group-wise effective tail exponents, lowers the aggregate effective tail exponent, and can drive the system across the condensation threshold $\gamma_{\rm c}=2$. These results identify volatility configuration as an additional control mechanism for condensation in nonequilibrium networked systems.
Figure 1: (a) Visualization of ER network with average degree $\langle k\rangle=4$. (b) Effective tail exponent $\hat{\gamma}$ in wealth distributions of the BM model for various network types against $\Lambda=2J/\beta^{2}$. The solid line shows the analytical baseline of the MF case. The other cases show numerical simulation results, of which symbols and error bars are the average and the standard deviation of 128 runs. The mean degrees are controlled by $\langle{k}\rangle=4.00\pm 0.027$. For all cases, $J=10^{-3}$ and $N=10^{4}$.
The remainder of this paper is organized as follows: In Sec. II, we introduce the heterogeneous BM model and analyze its properties in the homogeneous limit ($\Delta\beta\to 0$) for both the MF case and ER networks. Employing the SBM, we relate the assortativity $\mathcal{A}$ to the inter-block connection probability $q$. In Sec. III, we clarify the dependence of $\hat{\gamma}$ on $q$ and establish a configuration-induced wealth condensation transition. Finally, we summarize our findings and discuss the broader implications and possible directions for future study.
II Model
A power-law-tailed distribution is characterized by
$\rho(c)\sim c^{-\gamma}\quad\text{for large }c,$ (1)
where $\gamma$ is the power-law tail exponent. In particular, when $\gamma\leq 2$, the mean of an ideal tail sector diverges and the condensation transition occurs. The threshold $\gamma_{\rm c}=2$ therefore marks the condensation transition point, at which the inverse participation ratio (IPR), an order parameter [5], and the Gini index become singular.
We introduce binary volatility into the BM model. The stochastic differential equation (SDE) for the heterogeneous BM (HBM) model is given by
$dc_{i}=\frac{J}{\langle{k}\rangle}\sum_{j=1}^{N}a_{ij}(c_{j}-c_{i})\,dt+\beta_{i}c_{i}\,dW_{t,i},$ (2)
where $c_{i}$ denotes the normalized wealth [1] of node $i$, $t$ is time, $W_{t,i}$ is a Wiener process of $i$ at $t$, $\beta_{i}=\beta_{\pm}=\beta\pm\Delta\beta$ is the volatility of node $i$, $J$ is the strength of interaction between connected nodes, $\mathbf{a}$ is the adjacency matrix for a given network with the element $a_{ij}$, $\langle{k}\rangle$ is the mean degree, and $N$ is the total number of nodes. Throughout this paper, Eq. (2) is interpreted in the Itô interpretation. In contrast to the homogeneous model, the volatility configuration $\bm{\beta}=(\beta_{1},\dots,\beta_{N})$ for a given network plays the role of a quenched disorder and also influences a wealth distribution.
For a complete network, i.e., $a_{ij}=1-\delta_{ij}$, Eq. (2) reduces in the thermodynamic limit $N\to\infty$ to $dc_{i}=J(1-c_{i})\,dt+\beta_{i}c_{i}\,dW_{t}$. In the homogeneous limit $\Delta\beta\to 0$, the equation becomes the MF case of the BM model, whose stationary distribution is the inverse-gamma distribution with the tail exponent $\gamma=2+2J/\beta^{2}$. Even in the limit of $\Delta\beta\neq 0$, each volatility group still follows an effective MF result of the BM model, because the equation depends only on $c_{i}$ and $\beta_{i}$. The aggregate wealth distribution is therefore a mixture of two inverse-gamma distributions with exponents $\gamma_{\pm}=2+2J/\beta_{\pm}^{2}$. Hence, in the MF limit, the group-wise tail exponents remain unchanged despite the presence of heterogeneous volatility.
For ER networks, the power-law tail of $\rho(c)$ has also been numerically confirmed. Moreover, the effective tail exponent $\hat{\gamma}$ does not depend on the individual values of $J$ and $\beta^{2}$, but only on their ratio [6]. For convenience, we denote the control parameter as follows:
$\Lambda\equiv 2J/\beta^{2}.$
In the homogeneous limit of $\Delta\beta\to 0$, Fig. 1 shows the dependence of $\hat{\gamma}$ on $\Lambda$. Although the MF case of the BM model always yields $\gamma>2$ and therefore no condensation, the BM model on ER networks can exhibit $\hat{\gamma}\leq 2$ to undergo the wealth-condensation transition. For example, for $\langle{k}\rangle=4$, the transition occurs around $\Lambda\approx 0.4$. In detail, the effective tail exponent $\hat{\gamma}$ is estimated using a maximum-likelihood procedure together with a data-driven tail cutoff (Appendix B), and finite-size effects in tail fitting are assessed separately (Appendix C).
However, for $\Delta\beta\neq 0$, the exponent $\gamma_{\pm}$ differs substantially from the values obtained simply by substituting $\beta_{\pm}$ into the result of the homogeneous limit. The difference arises from local interactions between nodes with heterogeneous volatility, which is the key difference from the MF case. Since the value of $\gamma_{\pm}$ depends on how to allocate $\beta_{i}$ to nodes of a given network, we study this configuration effect on the HBM model.
[…]
Figure 3: (a) Aggregate effective tail exponent $\hat{\gamma}$ on the $(q,\Lambda_{2})$ plane at fixed $\Lambda_{1}=10^{-2}$. (b) Phase diagram of the condensed regime ($\hat{\gamma}\leq 2$) and the non-condensed regime ($\hat{\gamma}>2$). (c), (d) Group-wise effective tail exponents $\hat{\gamma}_{\pm}$. In all panels, the underlying network is an SBM with mean degree $\langle{k}\rangle=4.00\pm 0.027$, and we perform the simulations with $\Lambda_{1}=10^{-2}$, $J=10^{-3}$, and $N\approx 10^{4}$. The estimated exponents are averaged over 128 independent runs.
In order to characterize the configurational property that is relevant to the effective tail exponent, we consider the assortativity $\mathcal{A}$ of heterogeneous volatility, which measures how much similar volatility are connected in a given network:
$\mathcal{A}\equiv\frac{\mathrm{Cov}(\beta,\beta^{\prime})}{\sqrt{\mathrm{Var}(\beta)\mathrm{Var}(\beta^{\prime})}},$ (3)
where $\beta$ and $\beta^{\prime}$ denote the volatility of two connected nodes, respectively. In principle, $1\leq\mathcal{A}\leq 1$, although the attainable bounds depend on the details of the network structure [7]. When $N_{+}=N_{-}$, a random permutation of the binary volatility in an ER network typically yields $\mathcal{A}\approx 0$.
Although one may adjust $\mathcal{A}$ by repeatedly swapping pairs of nodes from an arbitrary initial configuration, this approach cannot exceed the structural limit imposed for a given network. It also makes it difficult to control other configurational properties independently, such as the volatility–degree correlation. The limitations can be naturally overcome by introducing a stochastic block model (SBM).
In this paper, we consider an SBM consisting of two equally sized blocks $a$ and $b$. Its block connection matrix is given by
$P_{ab}=\begin{bmatrix}p(1-q)&pq\\ pq&p(1-q)\end{bmatrix}$, (4)
where $p$ denotes the overall probability of connection between nodes (link) in the SBM, and $q$ is the probability of connection between blocks (inter-block connection). For a given mean degree $\langle{k}\rangle$, the corresponding ER network has a connection probability $p_{0}=\langle{k}\rangle/(N-1)$, while the SBM has $p\approx 2p_{0}$ in the large limit of $N$, so the mean degree is independent of $q$. In particular, when $q=1/2$, the SBM is equivalent to an ER network in which every pair of nodes is connected with probability $p/2$. The assignment of $\beta_{+}$ and $\beta_{-}$ to the two blocks yields
$\mathcal{A}=1-2q.$ (5)
Hence, the inter-block connection probability $q$ becomes a convenient control parameter to tune the configurational property $\mathcal{A}$ without significantly altering the overall structure of the networks; the near-equivalence of the ER and SBM degree statistics is summarized in Appendix A.
III Results
For the SBM, the wealth distribution exhibits nearly the same effective tail exponent $\hat{\gamma}$ as that of the ER network in the homogeneous limit $\Delta\beta\to 0$, since the degree distributions of the two networks are nearly identical. However, for $\Delta\beta\neq 0$, the effective tail exponents are highly dependent on $q$. At $q=0$, two blocks are completely separated ER networks. Consequently, each group has an effective tail exponent $\hat{\gamma}_{\pm}$ corresponding to $\Lambda_{1}=2J/\beta_{+}^{2}$ and $\Lambda_{2}=2J/\beta_{-}^{2}$, as in the homogeneous BM model [see Fig. 2(a)]. However, for $q>0$, local interactions between nodes with different volatility modify $\gamma_{\pm}$. Specifically, the difference between the two effective tail exponents decreases, producing a neutralization effect.
This trend is consistent with the form of the interaction term in Eq. (2), which redistributes wealth in proportion to the difference in normalized wealth between connected nodes. As $q$ increases, the nodes with different volatility interact more frequently, and the gap between the group-wise distributions and the effective tail exponent is reduced. Since this neutralization does not occur in the MF model, it originates from local interactions on the network.
In the homogeneous BM model, the effective tail exponent is determined solely by the control parameter $\Lambda$. In the heterogeneous BM model, by contrast, the configurational property of volatility acts as an additional control dimension that determines the effective tail exponent. To demonstrate this, we fix $\Lambda_{1}=2J/\beta_{+}^{2}$ and construct a phase diagram in the $(q,\Lambda_{2})$ plane, where $\Lambda_{2}=2J/\beta_{-}^{2}$. The effective aggregate tail exponent $\hat{\gamma}$ extracted from the overall distribution $\rho(c)$ varies not only with $\Lambda_{2}$ but also with $q$. Moreover, for appropriate values of $(q,\Lambda_{1},\Lambda_{2})$, the system undergoes a wealth-condensation transition across $\gamma_{\rm c}=2$, giving rise to a phase boundary in the $(q,\Lambda_{2})$ plane. As shown in Fig. 3, increasing $q$ lowers the effective tail exponent $\hat{\gamma}$, thus enhancing the overall inequality of the system.
The neutralization of $\hat{\gamma}_{\pm}$ does not affect the total distribution symmetrically. In all parameter regimes we tested, the group-wise distributions $\rho_{\pm}(c)$ show that the tail population is larger in the low-volatility group than in the high-volatility group. Consequently, the tail of the aggregate distribution $\rho(c)$ is governed primarily by the tail behavior of the $\beta_{-}$ group [see Fig. 3]. This is also reflected in the phase diagrams. Because the low-volatility group contributes much more strongly to the tail, the wealth shares of the two groups also differ substantially [see Fig. A3]. This gap in wealth share decreases as $q$ increases, indicating that inter-block connections hinder the net transfer of wealth from the high-volatility group to the low-volatility group.
IV Conclusion
We have investigated the role of heterogeneous volatility in a networked Bouchaud–Mézard model and shown that the configuration of volatility fundamentally alters the statistics of wealth distributions. In contrast to the homogeneous case, where the tail exponent is uniquely determined by the control parameter $\Lambda$, we find that the effective exponent $\hat{\gamma}$ depends additionally on the mixing structure of volatility, controlled here by the inter-block connection probability $q$. This configurational effect is absent in the MF limit and emerges only when heterogeneous volatility interacts locally on a sparse network.
A key mechanism underlying this behavior is the neutralization of the group-wise tail exponents through local exchange between nodes with different volatility. As volatility mixing increases, the disparity between subpopulations is reduced at the group level, yet the aggregate tail becomes heavier because the overall effective exponent decreases. In this way, wealth condensation can be induced not only by tuning a global parameter such as $\Lambda$ but also by changing the configurational correlation of noise intensity. The phase diagram on the $(q,\Lambda_{2})$ plane and the complementary robustness checks in Appendix C support this interpretation.
Our findings suggest a broader perspective on nonequilibrium systems with heterogeneous noise. Although earlier examples such as active matter [14, 4] and the random-field Ising model [12, 2] emphasize correlations in the noise itself, the present results show that correlations in noise amplitude (i.e., correlated volatility) can also control macroscopic phase behavior even when the driving noises remain mutually uncorrelated. In the context of wealth dynamics, this identifies volatility configuration as an additional control dimension for inequality formation and condensation.
Acknowledgements.
This research was supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) (KR) [NRF-RS-2026-25489888 (J.H., M.H.) and NRF-RS-2025-00514776 (J.H., H.J.)].
Appendix A Degree distributions for ER network and stochastic block model
In this appendix, we compare the network structures and degree distributions of the ER network and the stochastic block model (SBM). Let the total number of nodes be $N$ (assuming that it is even). In an ER network, the degree $k$ follows a binomial distribution,
$k\sim\mathrm{B}(N-1,p).$
Its mean is given by $\mathbb{E}[k]=(N-1)p$. Therefore, for a prescribed mean degree $k_{0}$, the connection probability of the ER network is
$p^{(\mathrm{ER})}=\frac{k_{0}}{N-1}.$
Figure A1: Degree distributions of a giant connected component (GCC) of the ER network and SBM with various $q$. The black dots are the theoretical probability mass function for the GCC of the ER network with $k_{0}=3.915$. The colored histograms show the degree distribution obtained by 100 independent runs of random network generations. For all cases, the mean degree is controlled as $\langle{k}\rangle=4.00\pm 0.027$.
[…]
For comparison, we consider a symmetric two-block SBM with equal block sizes, $N_{a}=N_{b}=N/2$. Its block connection matrix is given by Eq. (4). Let $X$ and $Y$ denote the numbers of intra-block and inter-block neighbors of a node, respectively. Then
$X\sim\mathrm{B}\left(\frac{N}{2}-1,\,p(1-q)\right),\qquad Y\sim\mathrm{B}\left(\frac{N}{2},\,pq\right),$
and these two random variables are independent. Hence, the degree of the SBM is
$k=X+Y.$
Its mean is
$\mathbb{E}[k]=\left(\frac{N}{2}-1\right)p(1-q)+\frac{N}{2}pq,$
which reduces to $\mathbb{E}[k]\simeq Np/2$. in the large limit of $N$. Therefore, for fixed $k_{0}$ and $N\to\infty$, the following relationship is established.
$p^{(\mathrm{SBM})}=2p^{(\mathrm{ER})}$ (6)
The limiting cases are straightforward. For $q=0$, all links are confined within blocks, so
$k=X+Y\sim\mathrm{B}\left(\frac{N}{2}-1,p\right),$
and the SBM reduces to two disconnected ER networks of size $N/2$, while preserving the same mean degree in the large limit of $N$. For $q=1/2$, the two binomial variables have the same success probability, and thus,
$k=X+Y\sim\mathrm{B}\left(N-1,\frac{p}{2}\right).$
Since $p=2p^{(\mathrm{ER})}$, the SBM is then exactly equivalent to an ER network of the same mean degree.
For $0<q<1/2$, the SBM is not strictly equivalent to the ER network, since $k$ is the sum of two binomial variables with different probabilities. Nevertheless, its mean and variance converge to those of the corresponding ER network as $N\to\infty$. Therefore, for sufficiently large $N$, the degree distribution and the overall network structure of the SBM are nearly identical to those of an ER network with a mean degree $k_{0}$.
Appendix B Details of numerical simulations
We first describe the numerical procedure used to estimate the effective tail exponent $\hat{\gamma}$. A linear fit to the double-logarithmic plot of $\rho(c)$ is highly sensitive to the histogram binning and does not provide estimates of $\hat{\gamma}$ consistent with the control parameter $\Lambda$. We therefore use the maximum-likelihood estimator (MLE) for the Pareto distribution,
$\hat{\gamma}=1+n\left[\sum_{i=1}^{n}\ln\left(\frac{c_{i}}{c_{\min}}\right)\right]^{-1},$ (7)
where $c_{i}$ denotes the normalized wealth satisfying $c_{i}>c_{\min}$, $c_{\min}$ is the lower cutoff for the tail, and $n$ is the number of samples. To estimate $c_{\min}$, we use the Kolmogorov-Smirnov (KS) statistic based on [8]. In summary, we obtain $c_{\min}$ from the goodness-of-fit with KS statistics and calculate $\hat{\gamma}$ from the MLE for the Pareto distribution for each run of the HBM simulation. Fig. 3 shows that averaged $\hat{\gamma}$ of 128 runs at $t=10^{5}$.
For network generation, we retain only the GCC, since disconnected nodes do not exchange wealth and would otherwise distort the stationary normalized wealth distribution. In an ER network, the degree distribution converges to the Poisson distribution for $N\to\infty$.
$p_{\mathrm{ER}}(k)=\frac{\lambda^{k}e^{-\lambda}}{k!},$
where $\lambda$ is the mean degree. The GCC degree distribution is
$p_{\mathrm{GCC}}(k)=\frac{\lambda^{k}e^{-\lambda}(1-u^{k})}{k!(1-u)}.$
Here $u$ satisfies the self-consistent equation $u=e^{-\lambda(1-u)}$. The mean degree of the GCC is
$\langle{k}\rangle_{\mathrm{GCC}}=\lambda-W_{0}(-\lambda e^{-\lambda}),$ (8)
where $W_{0}$ is a zeroth branch of the Lambert W function and $\langle{k}\rangle_{\rm GCC}$ is always larger than $\lambda$. Therefore, to obtain the target mean degree $\langle{k}\rangle_{\rm GCC}=4$ for the simulations, the ER network generation parameter must be chosen slightly below 4. So numerically, $\lambda\approx 3.915$ gives $\langle{k}\rangle_{\mathrm{GCC}}=4$.
The same preprocessing is applied to the SBM. After restricting to the GCC, the ER and SBM networks used in the simulations have nearly identical mean degree, degree distribution, and average path length. Accordingly, the corresponding SBMs yield almost the same wealth distributions and effective tail exponents $\hat{\gamma}$ in the homogeneous limit of $\Delta\beta\to 0$ [see Fig. 1(b)]. Since both ER network and SBM have almost identical degree distributions, their GCCs also have almost identical degree distributions regardless of $q$ [see Fig. A1].
Appendix C Complementary numerical simulations
C.1 Power-law test and size dependence
In this appendix, we examine the statistical test for power-law tails and their robustness across parameter regimes. Bouchaud and Mézard [6] reported that a random network with a mean degree $\langle{k}\rangle=4$ exhibits a power-law tail. Garlaschelli [10] discussed the relevance of log-normal behavior in sparsely connected clusters, whereas Souma et al[13] showed that, even at sparse and fixed mean degree, the shortcut effect makes small-world networks exhibit power-law tails. Motivated by these earlier observations, and going beyond visual inspection of double-logarithmic plots of normalized wealth distributions, we performed statistical comparisons based on the log-likelihood ratio test [15, 8].
[…]
These observations suggest that in the small-$\Lambda$ regime, the truncated log-normal fit may reflect finite-size rounding of the tail sector rather than the asymptotic form itself. In particular, as $N$ increases, the tail extends to larger $c$ and the fitted log-normal tail becomes progressively flatter, while the effective exponent $\hat{\gamma}$ remains approximately stable. This trend is consistent with a crossover to the power-law scaling in the thermodynamic limit of $N\to\infty$. We therefore interpret $\hat{\gamma}$ as an effective exponent against scaling, while avoiding the claim that finite-$N$ samples are uniquely characterized by a pure power-law over the entire fitted range. Within this interpretation, the cases with $\hat{\gamma}\leq 2$ are consistent with condensation in the thermodynamic limit.
C.2 Wealth distribution and asymmetric wealth shares in HBM model
Figure A3 shows the normalized wealth distributions and the group-wise wealth shares for the two volatility groups. A comparison of the group-wise wealth distributions indicates that the effective tail exponent of the overall distribution and wealth shares are primarily governed by the low-volatility group. As $q$ increases, the difference between the two groups in the tail sector becomes smaller, and accordingly the disparity in their wealth shares is also reduced.
When the effective tail exponent satisfies $\hat{\gamma}\ll 2$, the system is in the condensation regime, so even a very small difference in the tail populations is strongly amplified in the wealth shares. As a result, the decrease in the wealth-share difference with increasing $q$ remains negligible in this regime. By contrast, as $\hat{\gamma}$ increases, condensation-induced amplification weakens, and the difference in wealth share decreases more noticeably as $q$ increases.
Figure A3: (a) Double-logarithmic scaled plots of hroup-wise normalized wealth distributions $\rho_{\pm}(c)$ in HBM model on SBM against $c$ for $q=\{0.1,0.3,0.5\}$. (b) Corresponding group-wise wealth shares $\langle c_{\pm}\rangle$ against $q$. For all cases, $(\Lambda_{1},\Lambda_{2})=(0.1,1)$. All results are averaged over 128 independent runs, and error bars indicate standard deviations.
C.3 Modified numerical simulations
In this subsection, we briefly present additional numerical results for variants of the model that go beyond the binary-volatility setting and the symmetric two-block SBM considered in the main text.
We consider the following three cases.
  1. ()
    Continuous volatility within each block.
    Instead of assigning a single volatility value to each block, we draw volatility from a uniform distribution within each block:
    $\beta_{a}\sim\mathrm{Uniform}\!\left(\beta_{+}-\Delta\beta,\;\beta_{+}+\Delta\beta\right),$
    and
    $\beta_{b}\sim\mathrm{Uniform}\!\left(\beta_{-}-\Delta\beta,\;\beta_{-}+\Delta\beta\right).$
    Thus, the mean volatility in each block is still given by $\beta_{+}$ and $\beta_{-}$, respectively. This setup reduces to the one considered in the main text in the homogeneous limit of $\Delta\beta\to 0$. In the simulations, we set $\Delta\beta=|\beta_{+}-\beta_{-}|/2$ and examined the resulting group-wise wealth distributions.
  2. ()
    Asymmetric block sizes.
    We consider the case $N_{a}\neq N_{b}$. To control the mean degree $\langle{k}\rangle$ in the limit $N\to\infty$, we use the following block matrix:
    $P_{ab}=\begin{bmatrix}p_{a}(1-q)&\left(\frac{p_{a}+p_{b}}{2}\right)q\\ \left(\frac{p_{a}+p_{b}}{2}\right)q&p_{b}(1-q)\end{bmatrix}$, (9)
    where $p_{a}=\langle{k}\rangle/(N_{a}-1)$ and $p_{b}=\langle{k}\rangle/(N_{b}-1)$.
  3. ()
    Symmetric multi-block SBM.
    Finally, we consider a symmetric SBM with equal-sized blocks $M$. To control the mean degree $\langle{k}\rangle$ in the limit of $N\to\infty$, we use
    $P=\begin{bmatrix}p(1-q)&\frac{pq}{M-1}&\cdots&\frac{pq}{M-1}\\ \frac{pq}{M-1}&p(1-q)&\cdots&\frac{pq}{M-1}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{pq}{M-1}&\frac{pq}{M-1}&\cdots&p(1-q)\end{bmatrix}$, (10)
    where $p=\langle{k}\rangle/(N/M-1)$. The case $M=2$ reduces to the symmetric two-block setup studied in the main text. In the simulations for the multi-block case, we considered a symmetric four block SBM and assigned to the four block volatility values that are equally spaced on a logarithmic scale.
It is noted that when volatility is neither binary nor when block sizes are asymmetric, assortativity is no longer described by Eq. (5) of the main text. Nevertheless, in the symmetric cases, the degree–volatility correlation remains close to zero. The numerical results for these three cases are shown in Fig. A4. Consistent with the findings of the main text, the difference between the effective tail exponents in groups decreases as the inter-group connection probability $q$ increases.
Figure A4: From top to bottom panels, schematic illustrations of modified experimental setups (top), group-wise normalized wealth distributions $\rho_{\pm}(c)$ for $q=\{0.05,\,0.5\}$ (middle), and group-wise effective tail exponents $\hat{\gamma}_{\pm}$ against $q$ (bottom) are presented. From left to right, each columns represents corresponding numerical results, respectively, for (a) continuous volatility distributions within a block $(\beta_{a},\beta_{b})$, (b) an SBM with asymmetric block sizes ($N_{a}\neq N_{b}$), and (c) a symmetric multi-block SBM with $M=4$.

### 16 Apr 2026 v2

@@
 $dc_{i}=\frac{J}{\langle{k}\rangle}\sum_{j=1}^{N}a_{ij}(c_{j}-c_{i})\,dt+\beta_{i}c_{i}\,dW_{t,i},$ (2)
-where $c_{i}$ denotes the normalized wealth [1] of node $i$, $t$ is time, $W_{t,i}$ is a Wiener process of $i$ at $t$, $\beta_{i}=\beta_{\pm}=\beta\pm\Delta\beta$ is the volatility of node $i$, $J$ is the strength of interaction between connected nodes, $\mathbf{a}$ is the adjacency matrix for a given network with the element $a_{ij}$, $\langle{k}\rangle$ is the mean degree, and $N$ is the total number of nodes.
+where $c_{i}$ denotes the normalized wealth [1] of node $i$, $t$ is time, $W_{t,i}$ is a Wiener process of $i$ at $t$, $\beta_{i}=\beta_{\pm}=\beta\pm\Delta\beta$ is the volatility of node $i$, $J$ is the strength of interaction between connected nodes, $\mathbf{a}$ is the adjacency matrix for a given network with the element $a_{ij}$, $\langle{k}\rangle$ is the mean degree, and NN is the total number of nodes.
 Throughout this paper, Eq.
@@
 where $p$ denotes the overall probability of connection between nodes (link) in the SBM, and $q$ is the probability of connection between blocks (inter-block connection).
-For a given mean degree $\langle{k}\rangle$, the corresponding ER network has a connection probability $p_{0}=\langle{k}\rangle/(N-1)$, while the SBM has $p\approx 2p_{0}$ in the large limit of $N$, so the mean degree is independent of $q$.
+For a given mean degree $\langle{k}\rangle$, the corresponding ER network has a connection probability $p_{0}=\langle{k}\rangle/(N-1)$, while the SBM has $p\approx 2p_{0}$ in the large limit of NN, so the mean degree is independent of $q$.
 In particular, when $q=1/2$, the SBM is equivalent to an ER network in which every pair of nodes is connected with probability $p/2$.
@@
 In this appendix, we compare the network structures and degree distributions of the ER network and the stochastic block model (SBM).
-Let the total number of nodes be $N$ (assuming that it is even).
+Let the total number of nodes be NN (assuming that it is even).
 In an ER network, the degree $k$ follows a binomial distribution,
@@
 $\mathbb{E}[k]=\left(\frac{N}{2}-1\right)p(1-q)+\frac{N}{2}pq,$
-which reduces to $\mathbb{E}[k]\simeq Np/2$. in the large limit of $N$. Therefore, for fixed $k_{0}$ and $N\to\infty$, the following relationship is established.
+which reduces to $\mathbb{E}[k]\simeq Np/2$. in the large limit of NN. Therefore, for fixed $k_{0}$ and $N\to\infty$, the following relationship is established.
 $p^{(\mathrm{SBM})}=2p^{(\mathrm{ER})}$ (6)
@@
 $k=X+Y\sim\mathrm{B}\left(\frac{N}{2}-1,p\right),$
-and the SBM reduces to two disconnected ER networks of size $N/2$, while preserving the same mean degree in the large limit of $N$.
+and the SBM reduces to two disconnected ER networks of size $N/2$, while preserving the same mean degree in the large limit of NN.
 For $q=1/2$, the two binomial variables have the same success probability, and thus,
@@
 Nevertheless, its mean and variance converge to those of the corresponding ER network as $N\to\infty$.
-Therefore, for sufficiently large $N$, the degree distribution and the overall network structure of the SBM are nearly identical to those of an ER network with a mean degree $k_{0}$.
+Therefore, for sufficiently large NN, the degree distribution and the overall network structure of the SBM are nearly identical to those of an ER network with a mean degree $k_{0}$.
 Appendix B Details of numerical simulations
@@
 These observations suggest that in the small-$\Lambda$ regime, the truncated log-normal fit may reflect finite-size rounding of the tail sector rather than the asymptotic form itself.
-In particular, as $N$ increases, the tail extends to larger $c$ and the fitted log-normal tail becomes progressively flatter, while the effective exponent $\hat{\gamma}$ remains approximately stable.
+In particular, as NN increases, the tail extends to larger $c$ and the fitted log-normal tail becomes progressively flatter, while the effective exponent $\hat{\gamma}$ remains approximately stable.
 This trend is consistent with a crossover to the power-law scaling in the thermodynamic limit of $N\to\infty$.
-We therefore interpret $\hat{\gamma}$ as an effective exponent against scaling, while avoiding the claim that finite-$N$ samples are uniquely characterized by a pure power-law over the entire fitted range.
+We therefore interpret $\hat{\gamma}$ as an effective exponent against scaling, while avoiding the claim that finite-NN samples are uniquely characterized by a pure power-law over the entire fitted range.
 Within this interpretation, the cases with $\hat{\gamma}\leq 2$ are consistent with condensation in the thermodynamic limit.